OpenStudy (anonymous):

If angle E measures (7x -12)º, what is the measure of its complement?

OpenStudy (anonymous):

Complementary angles are angles that add up to 90. 7x - 12 + complementary angle = 90 90 - (7x -12) = complementary angle You can do the rest :)

OpenStudy (anonymous):

7x - 12 = 90 degrees because thats the complement

OpenStudy (anonymous):

7x = 102 solve for x x = 14.59

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Which one would it b (7x-192)º (102-7x)º (7x-102)º (192-7x)º

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

Btw i like physic's is incorrect

OpenStudy (anonymous):

you do not need to solve for x

OpenStudy (anonymous):

It is irrelevant to the question

OpenStudy (anonymous):

Yeahh i thought so when i saw my choices